OpenStudy (anonymous):

What is the vertex of the parabola whose equation is y = (x + 1)2 + 3?

OpenStudy (anonymous):

y = (x+1)^2+3 y-3 = (x+1)^2 compare with X^2= 4AY VERTEX=(-1,3)

OpenStudy (anonymous):

thanks